### v1

@@ 3.2 Monte Carlo sampling for 2D base polytopes
In this section, we apply the sampling algorithm to probe the set of all 2d base polytopes. For each 2D box, we take $N=100,000$ random samplings for each $n\in\{10,20,30\}$. We find that for $\mathcal{B}_{6}(\mathbb{P}^{2})$, the $GL(2,\mathbb{Z})$ symmetry factor is typically equal to 5. This follows because an arbitrary polytope inside $\mathcal{B}_{6}(\mathbb{P}^{2})$ can be transformed by an element in $D_{3}$ while still remaining in the box $\mathcal{B}_{6}(\mathbb{P}^{2})$, while a typical polytope is big enough compared to the box that there are no other transformations that keep it in the box. The $2\times 2$ matrix representation of these five $GL(2,\mathbb{Z})$ elements are
@@
We list the estimated total number of 2D base polytopes in Table 3, generated from each run with $N=100,000$ and a different $n$. As one can see, the total number of base polytopes peaks around the box $\mathcal{B}_{6}({\mathbb{F}_{9}})$. Note also that the estimated number of polytopes is fairly insensitive to the choice of $n$ in the range sampled.
@@
Now we combine the data from different boxes to form a comprehensive scan of 2d base polytopes. Note that it is possible that a particular base polytope $\Delta_{b,i}$ can fit into $t(\Delta_{b,i})\geq 1$ boxes, and we should use the formula (21) to estimate the total number of bases. To compute the number $t(\Delta_{b,i})$, we take a box $\mathcal{B}_{6}(B)$ and multiply a matrix $M\in GL(2,\mathbb{Z})$, on each vertex $v\in\Delta_{b,i}$, and form a new lattice polytope $\Delta_{b,i}^{\prime}=\{M\cdot v\}$. Then we explicitly check whether all the vertices of $\Delta_{b,i}^{\prime}$ are contained in the box $\mathcal{B}_{6}(B)$. If so, it means that $\Delta_{b,i}$ can be contained in the box $\mathcal{B}_{6}(B)$ after the $GL(2,\mathbb{Z})$ transformation $M$. We add up the number of boxes $\mathcal{B}_{6}(B)$ satisfying this condition, and use this as $t(\Delta_{b,i})$. Due to the fact that $GL(2,\mathbb{Z})$ is infinite, in practice we choose $M$ from the subset of $GL(2,\mathbb{Z})$ that are constructed as a product of no more than six generators of $GL(2,\mathbb{Z})$:
$\{\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\ ,\ \begin{pmatrix}0&1\\ 1&0\end{pmatrix}\ ,\ \begin{pmatrix}1&1\\ 0&1\end{pmatrix}\}\,$. (26)
@@
After combining the data in Table 3, we get the total number of 2d base polytopes
@@
With this data, we can confirm the Gaussian distribution of the $\log_{10}$ of weight factors discussed in Section 2.3. We plot in Figure 2 the number of sampling for different $\log_{10}(w_{i})$, with bin size of $0.1$ on the horizontal axis. We also show the distribution of the number of vertices $m$ in Figure 3, which also approximately fits into a Gaussian distribution.
@@ 3.3 Consistency checks
Using the same methods as in the last section to combine the different data from different boxes (21), we estimate the total number of base polytopes fitting in these boxes to be
@@
In Figure 6, we compare the exact number of base polytopes for each $h^{1,1}$ with the number inside each $k=5$ boxes. The combined curve is not a simple addition of the curves from different boxes, because as explained before, a base can generally fit into more than one boxes. It is interesting to note that the distribution of polytopes is not dominated by one or two of the boxes, even though some boxes are much larger than others, in the sense of containing more primitive lattice points.
@@ 4.1 Set of minimal $\mathcal{G}$-polytopes
We thus wish to enumerate the $\mathcal{G}$ polytopes with 4 and 5 vertices. As in Section 2.2, we transform each of these $\mathcal{G}$ polytope into the Hermite normal form.
@@
A natural question is: can we completely determine those of the above $\mathcal{B}_{6}(p,q)$ type. From the experience above with the $\mathcal{B}_{6}(p,q)$ distribution, it seems plausible that as long as we fix the first three vertices, the distribution of the fourth vertex will be fairly contiguous, or at most to have relatively small jumps between valid $\mathcal{G}$-polytopes. Keeping this idea in mind, we propose the following algorithm to search for the set of 3d minimal $\mathcal{G}$ polytopes (or at least the outline of the minimal box set). Here we take the case of 4 vertices as an example:
@@
    in the Hermite normal form 23 with $0<q$, $t\leq A$, $0\leq p<q$, $0\leq r,s<t$. Since we are considering $\mathcal{G}$ polytopes here, we do not require the vertices of the $\mathcal{G}$ polytope to be primitive. However, the $\mathcal{G}$ polytopes with non-primitive vertices are not minimal in general as one can imagine.
@@
We apply our algorithm on polytopes with $A=12,A_{4}=20,\Delta=6$. In other words, we enumerate $\mathcal{G}$ polytopes with the first three vertices containing only elements less than or equal to 12 and the fourth vertex with elements that are all negative but greater than 20-20, and explore around the minimal $\mathcal{G}$ polytopes already found. After modding out permutational redundancy, a total of 314 different $\mathcal{G}$ polytopes other than the $\mathcal{B}_{6}(p,q)$ type are found to be minimal. The explicit list of minimal $\mathcal{G}$ polytopes could be found in the attached file “3d-G-polytopes.txt” for the list of minimal $\mathcal{G}$-polytopes. Nonetheless, we do not provide an analytical, rigorous proof that there are all the minimal $\mathcal{G}$ polytopes.
Here are some comments on these results: The list of minimal $\mathcal{G}$ polytopes contain some examples on the boundary of the set we explored, so we believe it is probably not the complete set of minimal $\mathcal{G}$ polytopes. However, the number of minimal boxes decreases rapidly as the first three vertices move further from the origin. Minimal polytopes not in the $\mathcal{B}_{6}(p,q)$ style are a small subset (estimated at less than 500, and much smaller than 4000, the number of $\mathcal{B}_{6}(p,q)$ $\mathcal{G}$ polytopes) of all the minimal $\mathcal{G}$ polytopes. So we conclude that most of minimal polytopes are of the $\mathcal{B}_{6}(p,q)$ form, and in the following sections we will mainly focus on such examples. Moreover, we find that minimal $\mathcal{G}$ polytopes are concentrated in the pattern
@@
In total, we find 4454 inequivalent minimal $\mathcal{G}$ polytopes with 4 vertices, without proving the completeness.
As for the minimal $\mathcal{G}$ polytopes with 5 vertices, we have searched over $\mathcal{G}$ polytopes with 5 vertices and the maximal elements of the first three vertices less than 7. There are 98 $\mathcal{G}$ polytopes in total that are minimal within this set, of which 78 $\mathcal{G}$ polytopes of the form $\{(1,0,0),(0,1,0),(0,0,1),(0,-1,-p),(-1,0,-q)\}$ with $12\geq q\geq p\geq 1$. The other 20 minimal $\mathcal{G}$ polytopes are listed in table  5. Among them there is a special kind of minimal $\mathcal{G}$ polytope with a similar pattern
@@ 4.2 Monte Carlo in the largest box
There are $|R|=181203$ primitive rays in this base polytope. To generate compact base polytopes $\Delta_{b}$ in $\mathcal{B}_{6}(84,516)$, we have to fix the following two points
@@
We perform the Monte Carlo process with different numbers of sampled points $n=\{30,50,70,100,150\}$, and plot the estimated number of bases with respect to $h^{1,1}(B_{3})$ in Figure 10. One can see that when $n$ increases, the sampled bases generally have a larger $h^{1,1}(B_{3})$ as the plotted curves move to the top right. While for smaller $n$, the sampled bases have more variety and occupy a wider range of $h^{1,1}(B_{3})$.
Figure 10: The $\log_{10}$ of the estimated number of bases inside the largest box $\mathcal{B}_{6}(84,516)$, with two points $\{(-6,-6,1),(-6,37,-6)\}$ fixed, with respect to $h^{1,1}(B_{3})$. Different colors stand for data with different chosen $n$.
@@
Figure 11: The $\log_{10}$ of the estimated number of bases inside the largest box $\mathcal{B}_{6}(84,516)$, with two points $\{(-6,-6,1),(-6,37,-6)\}$ fixed, with respect to $h^{1,1}(B_{3})$. One adds up cases for all different $n$s to obtain a whole picture of how bases distribute with respect to $h^{1,1}(B_{3})$.
We verify the Gaussian distribution behaviours in Section 2.3. The distribution of weight factors acrossing samplings, for the combined data of different $n$s is shown in Figure 12, with a comparison to the optimal Gaussian fitting curve.
@@ 4.3 Monte Carlo in other boxes
As shown in Figure 9, the largest box $\mathcal{B}_{6}(84,516)$ lies at the bottom right of the figure. We consider the 1/4-, 1/2- and 3/4- positions of the diagonal line as examples of other boxes. The detailed information of the boxes are listed in table  6. Similarly, we fix the first two vertices to be (6,6,1),(6,37,6)(-6,-6,1)\ ,\ (-6,37,-6) to ensure it is possible to get an allowed base polytope, and randomly choose the remaining (n2)(n-2) vertices as in the previous section. The $h^{1,1}$ of different boxes ranges from $30000$ to $180000$. Finally, we collect all data with different sizes of boxes to form a more complete understanding of how bases with different sizes and shapes are distributed. The distribution of estimated number of bases with respect to $h^{1,1}$ from these four choices of boxes is shown in Figure 14.
@@ 4.4 Gauge group data
The different rigid gauge groups can be divided into two types. $SU(2),G_{2},F_{4},$ and $E_{8}$, are the most common gauge groups for a typical polytope. In Figures 15 to 18 we show the dependence of the average number of these gauge group factors for a single base. One can clearly see that the number of factors of these gauge groups increases similarly as $h^{1,1}$ grows, in a logarithm-like pattern. Notice that the common gauge groups have different slopes. The typical ratio of $SU(2):G_{2}:F_{4}:E_{8}$ is $50\%:33\%:13\%:4\%$. Note that the average numbers of each gauge group factor are different for different boxes even for the same $h$ value. This is due to our box Monte-Carlo strategy, leading to typical bases with different shapes and gauge groups in different boxes.
@@
Besides the four common gauge groups, the remaining five gauge groups $SU(3)$, $SO(7)$, $SO(8)$, $E_{6}$ and $E_{7}$ are relatively rare on a base. We use a logarithmic diagram to show the average number of rare gauge groups for different ranges of $h^{1,1}$. Generally, the average number of these gauge group factors per base is on the order of $10^{0}$ or less. Another interesting fact is that the sampling result from the largest box is qualitatively different from the smaller boxes. The sampled base polytopes in the largest box have much smaller numbers of rare gauge groups, and when $h^{1,1}>1.4\times 10^{5}$, we find a complete absence of these rare gauge groups in our data! With the results of the largest box, the rare gauge groups can be divided into two types. For gauge groups $G=SO(7)$ and $E_{7}$, even while the average number of such gauge group factors is small and gradually falls as $h^{1,1}$ grows, we generally see nonzero distributions in Figure 20 and 23. However, for $G=SU(3),SO(8),E_{6}$, the nonzero distributions are more sparse, and in most cases they would not appear in the bin of $h$; see Figure 19, 21 and 22.
@@ 4.5 Different base ensembles
Without a good handle on the cosmological measure problem, we really do not know which of these ensembles provides a better statistical measure on the F-theory landscape. Nonetheless, these provide different and useful windows on the variety of structures that can appear in this large landscape of consistent UV 4D quantum gravity theories. In the future a more physical measure could be computed via the quantum tunneling rate between different geometries, generalizing the weakly coupled IIB story in Gao:2025geq , possibly in an eternal inflation setup Hebecker:2020aqr .

### v2

@@ 3.2 Monte Carlo sampling for 2d base polytopes
In this section, we apply the sampling algorithm to probe the set of all 2d base polytopes [^6] The Mathematica code for the 2d Monte Carlo program is on https://doi.org/10.5281/zenodo.17761959 in the file 2d/MonteCarlo-2d.nb.. For each 2d box, we take $N=100,000$ random samplings for each $n\in\{10,20,30\}$. We find that for $\mathcal{B}_{6}(\mathbb{P}^{2})$, the $GL(2,\mathbb{Z})$ symmetry factor is typically equal to 5. This follows because an arbitrary polytope inside $\mathcal{B}_{6}(\mathbb{P}^{2})$ can be transformed by an element in $D_{3}$ while still remaining in the box $\mathcal{B}_{6}(\mathbb{P}^{2})$, while a typical polytope is big enough compared to the box that there are no other transformations that keep it in the box. The $2\times 2$ matrix representation of these five $GL(2,\mathbb{Z})$ elements are
@@
We list the estimated total number of 2d base polytopes in Table 3, generated from each run with $N=100,000$ and a different $n$. As one can see, the total number of base polytopes peaks around the box $\mathcal{B}_{6}({\mathbb{F}_{9}})$. Note also that the estimated number of polytopes is fairly insensitive to the choice of $n$ in the range sampled [^7] The Monte Carlo data are collected on https://doi.org/10.5281/zenodo.17761959 in the file 2d/2d-k=6.zip..
@@
Now we combine the data from different boxes to form a comprehensive scan of 2d base polytopes. Note that it is possible that a particular base polytope $\Delta_{b,i}$ can fit into $t(\Delta_{b,i})\geq 1$ boxes, and we should use the formula (22) to estimate the total number of bases. To compute the number $t(\Delta_{b,i})$, we take a box $\mathcal{B}_{6}(B)$ and multiply a matrix $M\in GL(2,\mathbb{Z})$ on each vertex $v\in\Delta_{b,i}$, and form a new lattice polytope $\Delta_{b,i}^{\prime}=\{M\cdot v\}$. Then we explicitly check whether all the vertices of $\Delta_{b,i}^{\prime}$ are contained in the box $\mathcal{B}_{6}(B)$. If so, it means that $\Delta_{b,i}$ can be contained in the box $\mathcal{B}_{6}(B)$ after the $GL(2,\mathbb{Z})$ transformation $M$. We add up the number of boxes $\mathcal{B}_{6}(B)$ satisfying this condition, and use this as $t(\Delta_{b,i})$. Due to the fact that $GL(2,\mathbb{Z})$ is infinite, in practice we choose $M$ from the subset of $GL(2,\mathbb{Z})$ that are constructed as a product of no more than six generators of $GL(2,\mathbb{Z})$:
$\left\{\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\ ,\ \begin{pmatrix}0&1\\ 1&0\end{pmatrix}\ ,\ \begin{pmatrix}1&1\\ 0&1\end{pmatrix}\right\}$,. (27)
@@
After combining the data in Table 3, we get the estimated total number of 2d base polytopes
@@
With this data, we can confirm the Gaussian distribution of the $\log_{10}$ of weight factors discussed in Section 2.3. We plot in Figure 2 the number of samplings for different $\log_{10}(w_{i})$, with bin size of $0.1$ on the horizontal axis. We also show the distribution of the number of vertices $m$ in Figure 3, which also approximately fits into a Gaussian distribution.
@@ 3.3 Consistency checks
Using the same methods as in the last section to combine the different data from different boxes (22), we estimate the total number of base polytopes fitting in these boxes to be [^9] The Monte Carlo data are collected on https://doi.org/10.5281/zenodo.17761959 in the file 2d/2d-k=5.zip.
@@
In Figure 6, we compare the exact number of base polytopes for each $h^{1,1}$ with the number inside each $k=5$ boxes. The combined curve is not a simple addition of the curves from different boxes, because as explained before, a base can generally fit into more than one box. It is interesting to note that the distribution of polytopes is not dominated by one or two of the boxes, even though some boxes are much larger than others, in the sense of containing more primitive lattice points.
@@ 4.1 Set of minimal $\mathcal{G}$-polytopes
We thus wish to enumerate the $\mathcal{G}$ polytopes with 4 and 5 vertices. As in Section 2.2, we transform each of these $\mathcal{G}$ polytopes into the Hermite normal form.
@@
A natural question is: can we completely determine those maximal boxes that are not of the above $\mathcal{B}_{6}(p,q)$ type. From the experience above with the $\mathcal{B}_{6}(p,q)$ distribution, it seems plausible that as long as we fix the first three vertices, the distribution of the fourth vertex will be fairly contiguous, or at most will have relatively small jumps between valid $\mathcal{G}$-polytopes. Keeping this idea in mind, we propose the following algorithm to search for the set of 3d minimal $\mathcal{G}$ polytopes (or at least the outline of the minimal box set). Here we take the case of 4 vertices as an example:
@@
    in the Hermite normal form (24) with $0<q$, $t\leq A$, $0\leq p<q$, $0\leq r,s<t$. Since we are considering $\mathcal{G}$ polytopes here, we do not require the vertices of the $\mathcal{G}$ polytope to be primitive. However, the $\mathcal{G}$ polytopes with non-primitive vertices are not minimal in general as one can imagine.
@@
We apply our algorithm on polytopes with $A=12$, $A_{4}=20$, $\Delta=6$. In other words, we enumerate $\mathcal{G}$ polytopes with the first three vertices containing only elements less than or equal to 12 and the fourth vertex with elements that are all negative but greater than 20-20, and explore around the minimal $\mathcal{G}$ polytopes already found. After modding out permutational redundancy, a total of 314 different $\mathcal{G}$ polytopes other than the $\mathcal{B}_{6}(p,q)$ type are found to be minimal. The explicit list of minimal $\mathcal{G}$ polytopes can be found on https://doi.org/10.5281/zenodo.17761959 in the file 3d/3d-G-polytopes.txt. Nonetheless, we do not provide an analytical, rigorous proof that these are all the minimal $\mathcal{G}$ polytopes.
Here are some comments on these results: The list of minimal $\mathcal{G}$ polytopes contain some examples on the boundary of the set we explored, so we believe it is probably not the complete set of minimal $\mathcal{G}$ polytopes. However, the number of minimal boxes decreases rapidly as the first three vertices move further from the origin. Minimal polytopes not in the $\mathcal{B}_{6}(p,q)$ style are a small subset (estimated at less than 500, and much smaller than 4000, the number of $\mathcal{B}_{6}(p,q)$ $\mathcal{G}$ polytopes) of all the minimal $\mathcal{G}$ polytopes. So we conclude that most of the minimal polytopes are of the $\mathcal{B}_{6}(p,q)$ form, and in the following sections we will mainly focus on such examples. Moreover, we find that minimal $\mathcal{G}$ polytopes are concentrated in the pattern
@@
In total, we find 4454 inequivalent minimal $\mathcal{G}$ polytopes with 4 vertices, without proving completeness.
As for the minimal $\mathcal{G}$ polytopes with 5 vertices, we have searched over $\mathcal{G}$ polytopes with 5 vertices and the maximal elements of the first three vertices less than 7. There are 98 $\mathcal{G}$ polytopes in total that are minimal within this set, of which 78 $\mathcal{G}$ polytopes are of the form $\{(1,0,0),(0,1,0),(0,0,1),(0,-1,-p),(-1,0,-q)\}$ with $12\geq q\geq p\geq 1$. The other 20 minimal $\mathcal{G}$ polytopes are listed in Table 5. Among them there is a special kind of minimal $\mathcal{G}$ polytope with a similar pattern
@@ 4.2 Monte Carlo in the largest box
There are $|R|=181203$ primitive rays in this base polytope. To generate compact base polytopes $\Delta_{b}$ in $\mathcal{B}=\mathcal{B}_{6}(84,516)$, we have to fix the following two points
@@
We perform the Monte Carlo process with different numbers of sampled points $n=\{30,50,70,100,150\}$, and plot the estimated number of bases with respect to $h^{1,1}(B_{3})$ in Figure 10. The data can be found on https://doi.org/10.5281/zenodo.17761959 in the file MonteCarlo3d-large.7z. One can see that when $n$ increases, the sampled bases generally have a larger $h^{1,1}(B_{3})$, as the plotted curves move to the top right. While for smaller $n$, the sampled bases have more variety and occupy a wider range of $h^{1,1}(B_{3})$.
Figure 10: The $\log_{10}$ of the estimated number of bases inside the largest box $\mathcal{B}_{6}(84,516)$, with two points $\{(-6,-6,1),(-6,37,-6)\}$ fixed, with respect to $h^{1,1}(B_{3})$. The full data can be accessed on https://doi.org/10.5281/zenodo.17761959 in the file MonteCarlo3d-large.7z. Different colors stand for data with different chosen $n$.
@@
Figure 11: The $\log_{10}$ of the estimated number of bases inside the largest box $\mathcal{B}_{6}(84,516)$, with two points $\{(-6,-6,1),(-6,37,-6)\}$ fixed, with respect to $h^{1,1}(B_{3})$. One adds up cases for all different $n$s to obtain a full picture of how bases are distributed with respect to $h^{1,1}(B_{3})$.
We verify the Gaussian distribution behaviours in Section 2.3. The distribution of weight factors across samplings, for the combined data of different $n$s, is shown in Figure 12, with a comparison to the optimal Gaussian fitting curve.
@@ 4.3 Monte Carlo in other boxes
As shown in Figure 9, the largest box $\mathcal{B}_{6}(84,516)$ lies at the bottom right of the figure. We consider the 1/4-, 1/2- and 3/4- positions of the diagonal line as examples of other boxes. The detailed information of the boxes are listed in table  6. Similarly, we fix the first two vertices to be (6,6,1),(6,37,6)(-6,-6,1)\ ,\ (-6,37,-6) to ensure it is possible to get an allowed base polytope, and randomly choose the remaining (n2)(n-2) vertices as in the previous section. The $h^{1,1}$ of different boxes ranges from $30000$ to $180000$. Finally, we collect all data with different sizes of boxes to form a more complete understanding of how bases with different sizes and shapes are distributed. One can get the data on https://doi.org/10.5281/zenodo.17761959 in the file MonteCarlo3d-medium.7z, MonteCarlo3d-small.7z and MonteCarlo-tiny.7z. The distribution of estimated number of bases with respect to $h^{1,1}$ from these four choices of boxes is shown in Figure 14.
@@ 4.4 Gauge group data
The different rigid gauge groups can be divided into two types. $SU(2),G_{2},F_{4},$ and $E_{8}$, are the most common gauge groups for a typical polytope. In Figures 15 to 18 we show the dependence of the average number of these gauge group factors for a single base. The data is given in 3d_gauge+weight files on https://doi.org/10.5281/zenodo.17761959. One can clearly see that the number of factors of these gauge groups increases similarly as $h^{1,1}$ grows, in a logarithm-like pattern. Notice that the common gauge groups have different slopes. The typical ratio of $SU(2):G_{2}:F_{4}:E_{8}$ is $50\%:33\%:13\%:4\%$. Note that the average numbers of each gauge group factor are different for different boxes even for the same $h$ value. This is due to our box Monte-Carlo strategy, leading to typical bases with different shapes and gauge groups in different boxes.
@@
Besides the four common gauge groups, the remaining five gauge groups $SU(3)$, $SO(7)$, $SO(8)$, $E_{6}$ and $E_{7}$ are relatively rare on a base. We use a logarithmic diagram to show the average number of rare gauge groups for different ranges of $h^{1,1}$. Generally, the average number of these gauge group factors per base is on the order of $10^{0}$ or less. Another interesting fact is that the sampling result from the largest box is qualitatively different from the smaller boxes. The sampled base polytopes in the largest box have much smaller numbers of rare gauge groups, and when $h^{1,1}>1.4\times 10^{5}$, we find a complete absence of these rare gauge groups in our data! With the results of the largest box, the rare gauge groups can be divided into two types. For gauge groups $G=SO(7)$ and $E_{7}$, even while the average number of such gauge group factors is small and gradually falls as $h^{1,1}$ grows, we generally see nonzero distributions in Figure 20 and 23. However, for $G=SU(3),SO(8),E_{6}$, the nonzero distributions are more sparse, and in most cases they would not appear in the bin of $h$; see Figures 19, 21 and 22.
@@ 4.5 Different base ensembles
Without a good handle on the cosmological measure problem, we really do not know which of these ensembles provides a better statistical measure on the F-theory landscape. Nonetheless, these provide different and useful windows on the variety of structures that can appear in this large landscape of consistent UV 4d quantum gravity theories, and there may be some utility to using each of these in studying various phenomenological aspects of the landscape (see, e.g., Cornell-axions). In the future a more physical measure could be computed via the quantum tunneling rate between different geometries, generalizing the weakly coupled IIB story in Gao:2025geq, possibly in an eternal inflation setup Hebecker:2020aqr.